OpenStudy (anonymous):

Let f(x)= 2x+3 and g(x)=-x^2+5 find (f*f)(x)

OpenStudy (unklerhaukus):

(f*f)(x)=\(f\circ f(x)\) ? (f*f)(x)=\(f\times f(x)\) ?

OpenStudy (anonymous):

Yes ! (2x+3)(2x+3) = 4x^2 +9 Am i correct ?

OpenStudy (unklerhaukus):

check your expansion

sam (.sam.):